OpenStudy (anonymous):

How many integers 1≤N≤1000 can be written both as the sum of 26 consecutive integers and as the sum of 13 consecutive integers?

OpenStudy (loser66):

exactly the same with the question above , take a look at http://openstudy.com/users/terenzreignz#/updates/51a32efee4b00f621d792b53

OpenStudy (raden):

shortly, try this lcm(13,26) = 26 floor(1000/26) = 38 so, it is 38 another way : Sum of 26 consecutive integers can be written as (-351+26a) 1<= -351+26a <= 1000 352 <= 26a <= 1351 13.5 <= a <= 51.96 Sum of 13 consecutive integers can be written as (-91+13b) 7.07 <= b <= 83.92 Accordingly informations above -351+26a = -91+13b 26a-13b = 260 2a-b = 20 b = 2a-20 a = (14,15,....51) satisfies it. So total 38 numbers

OpenStudy (anonymous):

Given that sin(2θ)=2/3, the value of sin^6θ+cos^6θ can be written as a/b with a and b coprime positive integers. what is a+b?

ganeshie8 (ganeshie8):

start by simplifying sin^6θ+cos^6θ

OpenStudy (anonymous):

What is the smallest 3-digit number which comprises of 3 distinct digits?

ganeshie8 (ganeshie8):

what do u think :)